# On the Convergence of Jacobian-Free Backpropagation for Optimal Control Problems with Implicit Hamiltonians

### v1

@@ 1 Introduction
$\dot{z}_{x}(t)$ $=-\nabla_{p}\mathcal{H}(t,z_{x}(t),p_{x}(t),u^{\star}(t)),,$ (3)
@@ 3.3 Training Problem Formulation
$u_{\theta}^{\star}$ $\in\operatorname*{arg\,max}_{u}\mathcal{H}(t,z,\nabla\phi_{\theta},u),$ (13c)
@@
The main computational challenge lies in evaluating and differentiating through the implicitly defined control $u_{\theta}^{\star}$. As discussed in Section 3.2, $u_{\theta}^{\star}$ is represented as the fixed point of an operator $T_{\theta}$ derived from the Hamiltonian optimality condition. For example, $T_{\theta}$ may be instantiated by an algorithmic update whose fixed points satisfy the first-order optimality condition om (13c), such as a gradient-based iteration, a proximal or operator-splitting scheme, or another monotone operator, while remaining differentiable with respect to the parameters $\theta$. However, this implicit representation introduces nontrivial computational costs during training, since differentiation must account for the dependence of the fixed point on $\theta$.
@@ 4.1 Notation and Essential Assumptions
The (true) gradient and JFB gradient approximation for a sample from the objective (13a)–(13c) as
@@
Assumption 4.4 is based on the original paper on JFB (Wu Fung et al., 2022), which requires conditioning restrictions on the JFB update $M_{\theta}$. For brevity, we denote $Var_{x}[v_{\theta,x}(t)]:=\mathbb{E}_{x}[\|v_{\theta,x}-E_{v}\|^{2}]$ and similarly for $Var_{x}[w_{\theta,x}(t)]:=\mathbb{E}_{x}[\|w_{\theta,x}-E_{w}\|^{2}]$.
@@
$\max(\text{Var}_{x}[v_{\theta,x}(t)],\text{Var}_{x}[w_{\theta,x}(t)])^{2}\leq\delta_{var}\|\mathbb{E}_{x}[M_{\theta}v_{\theta,x}]\|^{2},$ (20)
where $\lambda_{+}$ and $\lambda_{-}$ are uniform bounds on the largest and smallest eigenvalues of $M_{\theta}M_{\theta}^{\top}$ over $\theta,t,z,$ respectively.
Assumption 4.5 is possibly the most important assumption needed to prove convergence in expectation (and probability) to a critical point of (13a)–(13c). In words, this assumption controls how noisy the sample-wise integrands $v_{\theta,x}$ and $w_{\theta,x}$ are relative to their mean direction. This assumption is similar to variance bounds used standard proofs of stochastic gradient descent (SGD) (Bottou et al., 2018).
@@
Assumption 4.6 controls the magnitude of temporal deviations of the integrands from their time averages, which ensure that these deviations remain uniformly bounded. This condition captures a phenomenon specific to optimal control problems (the accumulation of gradient information over time) and does not arise in standard anslysis of SGD.
@@ 4.3 Convergence of SGD Using JFB as the Stochastic Gradient
Under the assumptions of Theorem  4.10, the JFB-based SGD iteration (22) satisfies
@@
Using Theorem  4.10, we can also prove convergence in probability to a critical point.
@@
Suppose the assumptions of Theorem  4.10 hold. For any $K\in\mathbb{N}$ let $j(K)\in\{0,1,...,K\}$ represent a random index chosen with probabilities proportional to $\{\alpha_{j}\}_{j=0}^{K}$. Then, $\{\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|\}_{j=0}^{K}\rightarrow 0$ as $K\rightarrow\infty$ in probability.
Finally, we note that it is possible to relax Assumption 4.5 and instead obtain convergence to a neighborhood of the stationary point instead C.4. See Section C for more details.
@@ 5.1 Theoretical Verification
While it is not practical to verify each remaining assumption leading to Lemma 4.9 individually, we instead demonstrate a consistent decrease in the training loss (right panel of Figure 2). This observed descent provides empirical evidence that the remaining assumptions are reasonable in practice and that the conclusion of Lemma 4.9 holds. Accordingly, the angle and descent checks presented here constitute the most meaningful empirical validation of the theory, as they directly assess whether the stochastic JFB updates act as descent directions.
@@ 5.2 Comparison in Low- and Moderate-Dimensional Settings
We compare several implicit deep learning approaches, including the proposed JFB method, CVXPYLayers (Agrawal et al., 2019), and traditional automatic differentiation (AD) (Paszke et al., 2019). To ensure a fair comparison, we focus on low- and moderate-dimensional problems, specifically, single-quadrotor and six-quadrotor control tasks, which remain tractable for all methods. For all experiments, learning rates were chosen to yield the best empirical performance. We used a fixed learning rate of $0.001$ for the single-quadrotor task, and a ReduceLROnPlateau scheduler with initial learning rate $0.01$, decay factor $0.5$, and patience $10$ for the six-quadrotor task.
@@
As shown in Figures 3 and 4, all methods exhibit comparable loss trajectories when measured against training epochs (top left panels). However, JFB converges substantially faster when performance is evaluated against wall-clock time (top right panels). This efficiency arises from avoiding both the storage of fixed-point iterations required by AD (bottom right) and the per-time-step linear solves required by CVXPYLayers. In addition, we visualize the resulting position trajectories for the quadrotor problem in Figure 5, which shows that JFB produces trajectories that are qualitatively comparable to those obtained using automatic differentiation and noticably closer to the target than those obtained using CVXPYLayers.
To provide a fairer comparison between AD and JFB, we also report loss versus work units, defined as the number of applications of $\partial T_{\theta}/\partial\theta$ (bottom left panels). We do not report this metric for CVXPYLayers, as its gradient computations are performed entirely as a black box; however, the reported runtimes indicate that it is likely less efficient than JFB but more efficient than AD. Finally, we note that the six-quadrotor setting represents the highest-dimensional problem that could be solved using AD before encountering memory limitations. Finally, Table 1 reports the final control objective values after training for both the single- and six-quadrotor experiments. While all methods achieve comparable final losses, JFB reaches these solutions with substantially faster runtimes and fewer work units its computational efficiency.
@@ 5.3 High-Dimensional Experiments
We next highlight the effectiveness of JFB on high-dimensional optimal control problems with implicit Hamiltonians. Traditional automatic differentiation (AD) is infeasible in these settings due to excessive memory consumption, as it requires storing every application of $T_{\theta}$ within the computational graph (see (Wu Fung et al., 2022) for a detailed discussion). CVXPYLayers is similarly infeasible for the consumption–savings and bicycle problems, since neither admits dynamics that are linear in the control variable $u$, and is prohibitively slow for the 100-quadrotor experiment. In contrast, JFB enables us to train these networks and approximate solutions for all three high-dimensional problems, as seen Figure 6 while also maintaining constant memory per epoch. Here, the 100 quadrotor problem is 1200-dimensional in the state and 400-dimensional in the controller. The 100 agent consumption problem is 100-dimensional in both state and controller. The 100 bikes are 400-dimensional in the state and 200-dimensional in the controller.
@@ 7 Conclusion
We develop the first convergence guarantees for Jacobian-Free Backpropagation (JFB) as a biased stochastic gradient method in optimal control problems with implicit Hamiltonians. Our results show that JFB converges to stationary points despite the absence of exact gradient information and enable scalable learning in high-dimensional control settings. Together, the theory and experiments establish JFB as a principled and practical approach for learning value function–based feedback controllers when closed-form Hamiltonians are unavailable. Future work will extend this framework to mean-field control/games settings (Lasry & Lions, 2007; Vidal et al., 2025; Agrawal et al., 2022; Laurière et al., 2022; Wang et al., 2025; Chow et al., 2022; Yang & Ichiba, 2023), where the scalability of our approach may prove useful. Code for this work will be provided upon publication.
@@ Appendix A Proof of Main Result
A.1 Proof of Lemma  4.7
Lemma  4.7: Under Assumptions 4.1-4.5,
@@
Thus, by the result of Lemma  4.7
@@
Thus, by the result of Lemma  4.8,  (52),  (53), and the fact that $\mathbb{E}_{\xi_{j}}[d_{\xi_{j}}^{JFB}(\theta_{j})]=\mathbb{E}_{x}[d_{x}^{JFB}(\theta_{j})]$, it holds that
@@
It can be derived that the RHS of  (54) is $\leq 0$ when $\alpha_{j}$ satisfies $0<\alpha_{j}\leq\frac{2\epsilon_{v}}{L_{J}(1-\gamma)^{2}}$. ∎
@@
Theorem 4.11: Under the assumptions of Theorem  4.10, the JFB-based SGD iteration (22) satisfies
@@
Let $K\in\mathbb{N}$ and $A_{K}=\sum_{k=0}^{K}\alpha_{k}$. Then, using linearity of expectation
$\mathbb{E}\left[\frac{1}{A_{K}}\sum_{j=0}^{K}\alpha_{k}\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|^{2}\right]=\frac{1}{A_{K}}\sum_{j=0}^{K}\mathbb{E}\left[\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|^{2}\right]$
@@
Corollary 4.12: Suppose the assumptions of Theorem  4.10 hold. For any $K\in\mathbb{N}$ let $j(K)\in\{0,1,...,K\}$ represent a random index chosen with probabilities proportional to $\{\alpha_{j}\}_{j=0}^{K}$. Then, $\{\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|\}_{j=0}^{K}\rightarrow 0$ as $K\rightarrow\infty$ in probability.
@@
$\leq\frac{1}{\epsilon^{2}}\mathbb{E}[\mathbb{E}_{j(K)}[\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j(K)})]]]$ (62)
By the proof of Theorem  4.10 we have $\lim_{K\rightarrow\infty}\mathbb{E}\left[\sum_{j=0}^{K-1}\alpha_{j}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}\right]<\infty$. Therefore, we must have $\lim_{j\rightarrow\infty}\mathbb{E}\left[\alpha_{j}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}\right]=0$. Thus, by  (62),
$\lim_{K\rightarrow\infty}P(\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j(K)})]\|\geq\epsilon)\leq\lim_{K\rightarrow\infty}\frac{1}{\epsilon^{2}}\mathbb{E}[\mathbb{E}_{j(K)}[\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j(K)})]]]=0$ (63)

### v2

@@ 1 Introduction
$\dot{z}_{x}(t)$ $=-\nabla_{p}\mathcal{H}(t,z_{x}(t),p_{x}(t),u^{\star}(t)),$ (3)
@@ 3.3 Training Problem Formulation
$u_{\theta}^{\star}$ $\in\operatorname*{arg\,max}_{u}\mathcal{H}(t,z_{x},\nabla\phi_{\theta},u),$ (13c)
@@
The main computational challenge lies in evaluating and differentiating through the implicitly defined control $u_{\theta}^{\star}$. As discussed in Section 3.2, $u_{\theta}^{\star}$ is represented as the fixed point of an operator $T_{\theta}$ derived from the Hamiltonian optimality condition. For example, $T_{\theta}$ may be instantiated by an algorithmic update whose fixed points satisfy the first-order optimality condition of (13c), such as a gradient-based iteration, a proximal or operator-splitting scheme, or another monotone operator, while remaining differentiable with respect to the parameters $\theta$. However, this implicit representation introduces nontrivial computational costs during training, since differentiation must account for the dependence of the fixed point on $\theta$.
@@ 4.1 Notation and Essential Assumptions
The (true) gradient and JFB gradient approximation for a sample from the objective (13a)–(13c) are defined as
@@
Assumption 4.4 is based on the original paper on JFB (Wu Fung et al., 2022), which requires conditioning restrictions on the JFB update $M_{\theta}$. For brevity, we denote $Var_{x}[v_{\theta,x}(t)]:=\mathbb{E}_{x}[\|v_{\theta,x}-E_{v}\|^{2}]$ and similarly for $Var_{x}[w_{\theta,x}(t)]:=\mathbb{E}_{x}[\|w_{\theta,x}-E_{w}\|^{2}]$, where $E_{v}=\mathbb{E}_{x}[v_{\theta,x}]$ and $E_{w}=\mathbb{E}_{x}[w_{\theta,x}]$.
@@
$\max\left(\sqrt{\text{Var}_{x}[v_{\theta,x}(t)]},\sqrt{\text{Var}_{x}[w_{\theta,x}(t)]}\right)^{2}\leq\delta_{var}\|\mathbb{E}_{x}[M_{\theta}v_{\theta,x}]\|^{2},$ (20)
where $\lambda_{+}$ and $\lambda_{-}$ are uniform bounds on the largest and smallest eigenvalues of $(M_{\theta}M_{\theta}^{\top})^{-1}$ over $\theta,t,z,$ respectively.
Assumption 4.5 is possibly the most important assumption needed to prove convergence in expectation (and probability) to a critical point of (13a)–(13c). In words, this assumption controls how noisy the sample-wise integrands $v_{\theta,x}$ and $w_{\theta,x}$ are relative to their mean direction. This assumption is similar to variance bounds used in standard proofs of stochastic gradient descent (SGD) (Bottou et al., 2018).
@@
Assumption 4.6 controls the magnitude of temporal deviations of the integrands from their time averages, which ensures that these deviations remain uniformly bounded. This condition captures a phenomenon specific to optimal control problems (the accumulation of gradient information over time) and does not arise in standard analysis of SGD.
Remark 4.7.
These assumptions are related to those in (Gelphman et al., 2025), but differ in several fundamental ways. In particular, the assumptions in (Gelphman et al., 2025) are imposed sample-wise, which are significantly stronger than those considered here, where all conditions are formulated in expectation with respect to $x$. Moreover, (Gelphman et al., 2025) assumes the existence of uniform lower bounds on $\|h_{\theta,x}\|$ and $\sigma_{\min}(M_{\theta})$, the smallest singular value of $M_{\theta}$. This distinction is crucial as such lower bounds generally fail to hold at stationary points, thereby precluding convergence guarantees under those assumptions. For this reason, we do not impose lower bounds on $\|h_{\theta,x}\|$ or $\sigma_{\min}(M_{\theta})$ in our analysis.
@@ 4.3 Convergence of SGD Using JFB as the Stochastic Gradient
Under the assumptions of Theorem  4.11, the JFB-based SGD iteration (22) satisfies
@@
Using Theorem  4.11, we can also prove convergence in probability to a critical point.
@@
Suppose the assumptions of Theorem  4.11 hold. For any $K\in\mathbb{N}$ let $j(K)\in\{0,1,...,K\}$ represent a random index chosen with probabilities proportional to $\{\alpha_{j}\}_{j=0}^{K}$. Then, $\{\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|\}_{j=0}^{K}\rightarrow 0$ as $K\rightarrow\infty$ in probability.
Finally, we note that it is possible to relax Assumption 4.5 and instead obtain convergence to a neighborhood of the stationary point instead. See Assumption  C.1 and Theorem  C.4 in Section C of the Appendix for more details on this weaker assumption. Moreover, under an assumption that is slightly stronger than Assumption  C.1 but significantly weaker than Assumption  4.5, to prove convergence in expectation to a critical point. See Assumption  D.1 and Theorem  D.4 in Section D of the Appendix for more information.
@@ 5.1 Theoretical Verification
While it is not practical to verify each remaining assumption leading to Lemma 4.10 individually, we instead demonstrate a consistent decrease in the training loss (right panel of Figure 2). This observed descent provides empirical evidence that the remaining assumptions are reasonable in practice and that the conclusion of Lemma 4.10 holds. Accordingly, the angle and descent checks presented here constitute the most meaningful empirical validation of the theory, as they directly assess whether the stochastic JFB updates act as descent directions. Moreover, it was not possible to perform the same numerical verification on larger experiments due to the excessive memory consumption of automatic differentiation needed to compute the exact gradient.
@@ 5.2 Comparison in Low- and Moderate-Dimensional Settings
We compare several implicit deep learning approaches, including the proposed JFB method, CVXPYLayers (Agrawal et al., 2019), and traditional automatic differentiation (AD) (Paszke et al., 2019). To ensure a fair comparison, we focus on low- and moderate-dimensional problems, specifically, single-quadrotor and six-quadrotor control tasks, which remain tractable for all methods. For all experiments, learning rates were chosen to yield the best empirical performance. Specifically, PyTorch’s ReduceLROnPlateau LR scheduler was used with initial learning rate $0.01$, decay factor $0.5$, and patience $10$.
@@
As shown in Figures 3 and 4, all methods exhibit comparable loss trajectories when measured against training epochs (top left panels). However, JFB converges substantially faster when performance is evaluated against wall-clock time (top right panels). This efficiency arises from avoiding both the storage of fixed-point iterations required by AD (bottom right) and the per-time-step linear solves required by CVXPYLayers. In addition, we visualize the resulting position trajectories for the quadrotor problem in Figure 5, which shows that JFB produces trajectories that are qualitatively comparable to those obtained using automatic differentiation and noticeably closer to the target than those obtained using CVXPYLayers.
To provide a fairer comparison between AD and JFB, we also report loss versus work units, defined as the number of applications of $\partial T_{\theta}/\partial\theta$ (bottom left panels). We do not report this metric for CVXPYLayers, as its gradient computations are performed entirely as a black box; however, the reported runtimes indicate that it is likely less efficient than JFB but more efficient than AD. Finally, we note that the six-quadrotor setting represents the highest-dimensional problem that could be solved using AD before encountering memory limitations. Finally, Table 1 reports the final control objective values after training for both the single- and six-quadrotor experiments. While all methods achieve comparable final losses, JFB reaches these solutions with substantially faster runtimes and fewer work units, which highlights its computational efficiency.
@@ 5.3 High-Dimensional Experiments
We next highlight the effectiveness of JFB on high-dimensional optimal control problems with implicit Hamiltonians. Traditional automatic differentiation (AD) is infeasible in these settings due to excessive memory consumption, as it requires storing every application of $T_{\theta}$ within the computational graph (see (Wu Fung et al., 2022) for a detailed discussion). CVXPYLayers is similarly infeasible for the consumption–savings and bicycle problems, since neither admits dynamics that are linear in the control variable $u$, and is prohibitively slow for the 100-quadrotor experiment. In contrast, JFB enables us to train these networks and approximate solutions for all three high-dimensional problems, as seen in Figure 6 while also maintaining constant memory per epoch. Here, the 100 quadrotor problem is 1200-dimensional in the state and 400-dimensional in the controller. The 100 agent consumption problem is 100-dimensional in both state and controller. The 100 bikes are 400-dimensional in the state and 200-dimensional in the controller.
@@ 7 Conclusion
We develop the first convergence guarantees for Jacobian-Free Backpropagation (JFB) as a biased stochastic gradient method in optimal control problems with implicit Hamiltonians. Our results show that JFB converges to stationary points despite the absence of exact gradient information and enable scalable learning in high-dimensional control settings. Together, the theory and experiments establish JFB as a principled and practical approach for learning value function–based feedback controllers when closed-form Hamiltonians are unavailable. Future work will extend this framework to mean-field control/games settings (Lasry & Lions, 2007; Vidal et al., 2025; Agrawal et al., 2022; Laurière et al., 2022; Wang et al., 2025; Chow et al., 2022; Yang & Ichiba, 2023; Yang, 2026), where the scalability of our approach may prove useful. Code for this work will be provided upon publication.
8 Acknowledgements
SWF and EG were partially funded by NSF Award 2309810. SO was partially funded by DARPA under grant HR00112590074, NSF under grant 2208272 and 1554564, AFOSR under MURI grant N00014-20-1-278, and by ARO under grant W911NF-24-1-015. SWF and EG were partially funded by NSF Award 2309810. SO was partially funded by DARPA under grant HR00112590074, NSF under grant 2208272 and 1554564, AFOSR under MURI grant N00014-20-1-278, and by ARO under grant W911NF-24-1-015.
@@ Appendix A Proof of Main Result
A.1 Proof of Lemma  4.8
Lemma  4.8: Under Assumptions 4.1-4.5,
@@
Thus, by the result of Lemma  4.8
@@
Thus, by the result of Lemma  4.9,  (52),  (53), and the fact that $\mathbb{E}_{\xi_{j}}[d_{\xi_{j}}^{JFB}(\theta_{j})]=\mathbb{E}_{x}[d_{x}^{JFB}(\theta_{j})]$, it holds that
@@
For the RHS of  (54) to be $\leq 0$, it must be true that, with $\alpha_{j}>0$
$\frac{\alpha_{j}^{2}L_{J}B_{max}^{2}T^{2}}{2\beta}$ $\leq\alpha_{j}\epsilon_{v}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}$ (55)
$\alpha_{j}$ $\leq\frac{2\beta\epsilon_{v}}{L_{J}B_{max}^{2}T^{2}}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}$ (56)
The following upper bound, uniform in $j$, for $\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}$ can be derived in a very similar manner to that of $\mathbb{E}_{x}[\|d_{x}^{JFB}(\theta_{j})\|^{2}]$ above,
$\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}\leq\frac{B_{max}^{2}T^{2}}{\beta(1-\gamma)^{2}}$ (57)
Thus, for the RHS of (54) to be $\leq 0$, it is necessary that
$\alpha_{j}\leq\frac{2\epsilon_{v}}{L_{J}(1-\gamma)^{2}}$ (58)
@@
Theorem 4.12: Under the assumptions of Theorem  4.11, the JFB-based SGD iteration (22) satisfies
@@
Let $K\in\mathbb{N}$ and $A_{K}=\sum_{j=0}^{K-1}\alpha_{j}$. Then, using linearity of expectation
$\mathbb{E}\left[\frac{1}{A_{K}}\sum_{j=0}^{K-1}\alpha_{j}\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|^{2}\right]=\frac{1}{A_{K}}\sum_{j=0}^{K-1}\alpha_{j}\mathbb{E}\left[\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|^{2}\right]$
@@
Corollary 4.13: Suppose the assumptions of Theorem  4.11 hold. For any $K\in\mathbb{N}$ let $j(K)\in\{0,1,...,K\}$ represent a random index chosen with probabilities proportional to $\{\alpha_{j}\}_{j=0}^{K}$. Then, $\{\left\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\right\|\}_{j=0}^{K}\rightarrow 0$ as $K\rightarrow\infty$ in probability.
@@
$\leq\frac{1}{\epsilon^{2}}\mathbb{E}[\mathbb{E}_{j(K)}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j(K)})]\|^{2}]]$ (66)
By the proof of Theorem  4.11 we have $\lim_{K\rightarrow\infty}\mathbb{E}\left[\sum_{j=0}^{K-1}\alpha_{j}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}\right]<\infty$. Therefore, we must have $\lim_{j\rightarrow\infty}\mathbb{E}\left[\alpha_{j}\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j})]\|^{2}\right]=0$. Thus, by  (66),
$\lim_{K\rightarrow\infty}P(\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j(K)})]\|\geq\epsilon)\leq\lim_{K\rightarrow\infty}\frac{1}{\epsilon^{2}}\mathbb{E}[\mathbb{E}_{j(K)}[\|\mathbb{E}_{x}[\nabla_{\theta}J_{x}(\theta_{j(K)})]\|^{2}]]=0$ (67)